# supplementary

### v2

@@ 2.3. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ can be analytically derived when the prior is non-informative on $\beta$. By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $(1-\rho)B^{-1}=U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$. Here, $U$ does not depend on $\rho$ and $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}$ is diagonal with elements $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}$and we immediately obtain $\lim_{\rho\rightarrow 0^{+}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=0_{n\times n}$. Next, $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$, where $\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)_{ii}=d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}$ which is 0 for any $\rho\in(0,1)$ if $d_{i}=0$, and converges to 0 as $\rho\rightarrow 0^{+}$ if $d_{i}>0$. Thus, $\lim_{\rho\rightarrow 0^{+}}\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)=0_{n\times n}$. These imply that if the prior is non-informative on $\beta$, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
@@
Turning to $\rho\rightarrow 1^{-}$, we note that if $d_{i}=0$, $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\rho$, which converges to $1$ as $\rho\rightarrow 1^{-}$. If $d_{i}>0$, then $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}=0$ and $\lim_{\rho\rightarrow 1^{-}}d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}=1$. Thus, $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$, and $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=UD^{\star}U^{-1}e$, where $e=(I-H)y$. Second order moments have limits given by $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}V$, and $\lim_{\rho\rightarrow 1^{-}}\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Furthermore, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis). Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$.
@@ 3.1. Connection to Classical Hypothesis Testing
Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test a point null hypotheses $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$ using the classical linear model, $Y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is a fixed but unknown $p\times 1$ vector (see, e.g., seberHypothesisTesting2003). For each $c_{k}$, this goal is accomplished by computing the test statistic $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $\widehat{\sigma}^{2}=\frac{1}{n-p}\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates. The classical approach computes a p-value $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$ where $T\sim t_{n-p}$ and accounts for multiple testing by computing adjusted p-values. If the adjusted p-value is below a pre-specified significance threshold, then $H^{(k)}_{0}$ is rejected.
In the Bayesian framework, which we adopt here, we maintain the continuous conjugate prior specified by (2) and recast the decision problem to consider the posterior probabilities of absolute standardized linear combinations among elements of $\beta$ being greater than a threshold $\epsilon$. For all $k=1,\ldots,K$ and any $\epsilon>0$, we define
@@
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations and represents the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from zero. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (9) using
@@
for any given $\epsilon$. Using a conjugate prior, we efficiently obtain exact posterior samples of $\beta\,|\,y$ using (3). Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which may, for example, represent mean outcomes for different treatment groups in an ANOVA.
Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (9) have opposite interpretation: a small p-value, say $p_{k}$, suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0. The posterior distribution of $\beta$ with a flat prior is connected to the generalized least squares estimate $\widehat{\beta}$ in the Bayesian linear regression model. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities are perfectly correlated.
@@
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$, and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$ for any given $\epsilon>0$.
@@
See Appendix LABEL:append:prop1_proof.
For the linear model, the Bayesian approach of ranking $\epsilon$-difference probabilities can replicate the rankings of the classical p-values by using a flat $\beta$ prior. Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
@@ 3.2. Extension to Spatial Random Effects
We next consider the model in (4), where each observation corresponds to one areal region. We treat $V_{\phi}$ as a fixed random effects variance structure, whereas $\rho$, which controls the proportion of the total error variance attributed to the structured component, can either be a fixed hyperparameter or a random variable with a prior density. Generalizing the idea of $\epsilon$-difference probabilities in the fixed effects model, we now conduct posterior inference on random variables of the form $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$, for any given $\epsilon>0$. For $k=1\ldots,K$, we define the conditional posterior probability
@@
where $c_{k}\in\mathbb{R}^{n+p}$ is typically a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (7).
In the context of detecting spatial disparities, consider a set of $K$ pairs of neighboring regions denoted as $L=\{(i,j):i<j,i\sim j\}$ where $i\sim j$ means that regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then we denote the corresponding $\epsilon$-difference probability as $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. The set $\left\{h_{ij}(\epsilon;\rho)\right\}_{(i,j)\in L}$ then represents the $\epsilon$-difference probabilities of all boundaries under consideration.
According to the equivalence between the model given by (4) and the augmented linear system discussed in Section 2.2, similar insight to the proof of Theorem 1 reveals that the rankings of the conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.

### v3

@@ 2.2. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$. By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$. Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$. Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
@@
We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$. Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis). Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$. These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).
@@ 3.1. Connection to Classical Hypothesis Testing
We first briefly address testing for contrasts in Bayesian linear regression without spatial effects, i.e., $\gamma=0$ in (1). Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003). We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates. We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate posterior probabilities for the absolute standardized linear combinations among elements of $\beta$ to exceed a threshold $\epsilon$. Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
@@
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 00. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (6) using
@@
for any given $\epsilon$. Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We draw exact posterior samples from the Normal-Gamma family. Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean outcomes for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 00, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 00. In fact, Bayesian inference with a flat prior on $\beta$ amounts to using the classical generalized least squares estimate $\widehat{\beta}$. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities yield identical decisions on contrasts in the following sense.
@@
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$, and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
@@
See LABEL:append:prop1_proof.
Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
@@ 3.2. Extension to Spatial Random Effects
We consider (1), where each observation corresponds to one region. We let $V_{\phi}$ be known, whereas $\rho$ can either be a fixed hyperparameter or a random variable with a prior density. Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we conduct posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$. For $k=1\ldots,K$, we define
@@
where $c_{k}\in\mathbb{R}^{n+p}$ is a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (4).
In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the rankings of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.